OpenStudy (anonymous):

Integrate dx/(1+sqrt(1-x^2))

OpenStudy (anonymous):

Do I need to rationalize it?

OpenStudy (shubhamsrg):

thats what I was going to write .

OpenStudy (anonymous):

multiply it by 1+sqrt(1-x^2)/1+sqrt(1-x^2)?

OpenStudy (shubhamsrg):

nop, in rationalizing, you multiply divide by the conjugate the conjugate here is 1-sqrt(1-x^2)

OpenStudy (anonymous):

I thought I was typing "-" lol

OpenStudy (anonymous):

So I got integral of 1- sqrt(1-x^2)/x^2 dx

OpenStudy (anonymous):

the original problem is integrate dx/1+ sqrt(1-x^2). Can I sue that identity there?

OpenStudy (jhannybean):

no.

OpenStudy (anonymous):

So I really need to use the rationalization

OpenStudy (jhannybean):

Yeah lol,i read the problem wrong. -_- haha sorry:(

OpenStudy (anonymous):

from rationalizing it, I got integral of 1- sqrt(1-x^2)/x^2 dx

OpenStudy (anonymous):

oh lol something's wrong i guess

OpenStudy (shubhamsrg):

now you can split the integral like this : integral (1/x^2) - integral (sqrt(1-x^2)/x^2) first part is easy, for second part, perhaps integration by parts might help ?

OpenStudy (anonymous):

I'm now at that. Thanks.

OpenStudy (shubhamsrg):

oh wait for 2nd part, do this substitute x=sint this'll make things much easier

OpenStudy (jhannybean):

\[\large \frac{1}{1+\sqrt{1-x^2}}\cdot \frac{1-\sqrt{1-x^2}}{1-\sqrt{1-x^2}}\]\[\large \frac{1-\sqrt{1-x^2}}{1-(1-x^2)} \]\[\large \frac{1-\sqrt{1-x^2}}{x^2}\]

OpenStudy (anonymous):

I've splitted it now like shubhamsrg said

OpenStudy (jhannybean):

hmm... That's what i got after multiplying by the conjugate.

OpenStudy (jhannybean):

Ok.

OpenStudy (jhannybean):

\[\large \int \left( \frac{1-\sqrt{1-x^2}}{x^2}\right)dx \]\[\large \int \frac{1}{x^2}dx - \int \left(\frac{\sqrt{1-x^2}}{x^2}\right)dx \]

OpenStudy (anonymous):

What should I do with the second part lol

OpenStudy (shubhamsrg):

as I said above, substitute x= sint

OpenStudy (anonymous):

didn't see your second reply, sorry

OpenStudy (anonymous):

that t is just a variable?

OpenStudy (shubhamsrg):

yes

OpenStudy (jhannybean):

\[\large x= \sin(t) \ , \ dx = \cos(t)dt \]\[\large \int \left(\frac{\sqrt{1-\sin^2(t)}}{\sin^2(t)} \right)dt\]

OpenStudy (anonymous):

yay

OpenStudy (shubhamsrg):

nop a slight mistake there

OpenStudy (anonymous):

uhm, where?

OpenStudy (shubhamsrg):

it should be cost dt

OpenStudy (anonymous):

the dt

OpenStudy (anonymous):

yeah. saw that now

OpenStudy (jhannybean):

yeah forgot to include that.

OpenStudy (jhannybean):

sorry -_-

OpenStudy (anonymous):

so it will be integral of cot^2 u du

OpenStudy (jhannybean):

Yes.

OpenStudy (shubhamsrg):

yes, thats right rest should be easier

OpenStudy (jhannybean):

and \[\large t= \sin^{-1}(t)\]

OpenStudy (shubhamsrg):

note that cot^2 u = cosec^2 u -1 and also derivative of cot u is -cosec^2 u

OpenStudy (jhannybean):

\[\large \int\limits \left(\frac{\sqrt{1-\sin^2(t)}}{\sin^2(t)} \right)(\cos(t))dt\]\[\large \int\limits \left(\frac{\cos^2(t)}{\sin^2(t)} \right)dt\]\[\large \int\limits \left(\cot^2(t) \right)dt\]\[\large -\csc^2(t) +c\]\[\large -\csc^2(\sin^{-1}(t))+c \]

OpenStudy (jhannybean):

BLEH noooo!!!

OpenStudy (shubhamsrg):

.-.

OpenStudy (jhannybean):

NOOO!!! hold on lol.

OpenStudy (anonymous):

I'm confused lol After I integrate cot^2 u du, how do I substitute it back to x?

OpenStudy (jhannybean):

\[\large \int\limits \cot^2(t)dt\]\[\large \int\limits(-1 + \csc^2(t))dt\]\[\large -t -\cot(t)+c\]

OpenStudy (jhannybean):

That's the tricky part, subbing back.

OpenStudy (anonymous):

I'll be using arc tan something?

OpenStudy (jhannybean):

Create a triangle, that's what i would do.

OpenStudy (jhannybean):



OpenStudy (anonymous):

my final answer is -1/x + sin^-1(x) + sqrt(1-x^2)/x + C

OpenStudy (jhannybean):

GOODJOB!!!!!!!

OpenStudy (jhannybean):

I'm writing it up using equation editor.

OpenStudy (anonymous):

Thank you both. anyway, I can't seem to verify my newly made account here , I've sent the verification link to my email may times

OpenStudy (anonymous):

Can I also ask questions in Physics here?

OpenStudy (jhannybean):

\[\large x= \sin(t) \ \therefore \ t = \sin^{-1}(x)\]\[\large \cot(t) = \frac{\cos(t)}{\sin(t)}=\frac{\sqrt{1-x^2}}{1}\cdot \frac{1}{x}= \frac{\sqrt{1-x^2}}{x} \]\[ -t-\cot(t)+c =-\sin^{-1}(x)-\frac{\sqrt{1-x^2}}{x} +c = -\frac{x\sin^{-1}(x)-\sqrt{1-x^2}}{x}+c \]

OpenStudy (jhannybean):

I'm lagging hardcore.

OpenStudy (jhannybean):

And there is a physics section on OS,too. just click "find more subjects"

OpenStudy (anonymous):

Okay. Thank you! I'll close this now :)

OpenStudy (jhannybean):

okie dokie.

OpenStudy (irishboy123):

shouldn't you just start with the trig substitution right at the outset? surely much easier to do?

OpenStudy (jhannybean):

Yes you are right, lol. Or use hyperbolic subs....

OpenStudy (irishboy123):

@Jhannybean hypernice!